OpenStudy (anonymous):

If f(x) = x + 4, find f(3). 1 7 −1 12

OpenStudy (anonymous):

3+4 = ?

OpenStudy (anonymous):

7

OpenStudy (anonymous):

f(3) = 3 + 4 = ??? just replace the x with a 3 and simplify...:)

OpenStudy (maheshmeghwal9):